Contents

Preface ix János Kollár

Fundamental Groups of Smooth Projective Varieties 1 Donu Arapura

Vector Bundles on Curves and Generalized Theta Functions: Recent Results and Open Problems 17

Arnaud Beauville

Recent Results in Higher Dimensional Birational Geometry 35 Alessio Corti

The Schottky Problem: An Update 57 Olivier Debarre

Spectral Covers 65 Ron Donagi

Adjoint Linear Systems 87 Lawrence Ein

Torelli Groups and Geometry of Moduli Spaces of Curves 97 Richard M. Hain

Vector Bundles and Brill–Noether Theory 145 Shigeru Mukai